# ANVIL: Anomaly-based Vulnerability Identification without Labelled Training Data

### v1

@@ Abstract
Supervised learning-based software vulnerability detectors often fall short due to the inadequate availability of labelled training data. In contrast, Large Language Models (LLMs) such as GPT-4, are not trained on labelled data, but when prompted to detect vulnerabilities, LLM prediction accuracy is only marginally better than random guessing. In this paper, we explore a different approach by reframing vulnerability detection as one of anomaly detection. Since the vast majority of code does not contain vulnerabilities and LLMs are trained on massive amounts of such code, vulnerable code can be viewed as an anomaly from the LLM’s predicted code distribution, freeing the model from the need for labelled data to provide a learnable representation of vulnerable code. Leveraging this perspective, we demonstrate that LLMs trained for code generation exhibit a significant gap in prediction accuracy when prompted to reconstruct vulnerable versus non-vulnerable code.
Using this insight, we implement ANVIL, a detector that identifies software vulnerabilities at line-level granularity. Our experiments explore the discriminating power of different anomaly scoring methods, as well as the sensitivity of ANVIL to context size. We also study the effectiveness of ANVIL on various LLM families, and conduct leakage experiments on vulnerabilities that were discovered after the knowledge cutoff of our evaluated LLMs. On a collection of vulnerabilities from the Magma benchmark, ANVIL outperforms state-of-the-art line-level vulnerability detectors, LineVul and LineVD, which have been trained with labelled data, despite ANVIL having never been trained with labelled vulnerabilities. Specifically, our approach achieves 1.62×1.62$\times$1.62 × to 2.18×2.18$\times$2.18 × better Top-5 accuracies and 1.02×1.02$\times$1.02 × to 1.29×1.29$\times$1.29 × times better ROC scores on line-level vulnerability detection tasks.
@@ I Introduction
Probabilistic approaches to static vulnerability detection have garnered considerable research attention due to the promise of automatically detecting common bug patterns by training a model on historical data. Confidence in statistical modelling has only grown with the recent advances in large language models (LLMs). One of the significant advantages of LLMs is that they can be trained in a self-supervised fashion without labelled training data. This has enabled them to be trained on very large training sets. This approach has led to impressive and surprising results in code understanding and code generation [1, 2].
@@
Unfortunately, the same cannot be said about their ability to detect vulnerabilities. As shown in a recent study [3], even with cutting edge LLMs like GPT-4, vulnerability prediction accuracy on real world code is little better than a coin-flip. Current learning and LLM-based vulnerability detectors [4, 5, 6, 7, 8, 9, 10, 11] suffer from the same lack of a large and well-labelled set of representative vulnerabilities. Attempts to compensate for the dearth of labelled data by selecting relevant code features (e.g. the granularity of the code context, commit history, control-flow, data-flow, etc.) and converting these features into a learnable representation (e.g. a vector, a graph, etc.) have only yielded modest improvements. Similarly, recent attempts to generate new vulnerable training corpora [12, 13, 14] have not improved real-world performance significantly, as synthetically generated vulnerabilities still cannot capture the huge variety of ways that vulnerabilities can manifest.
@@
LLMs are effective because they can learn a general feature space from a large training set. This does not lend itself well to understanding vulnerable code patterns due to a fundamentally imbalanced corpora. Out of the huge amounts of unlabelled code that LLMs are trained on, only a miniscule portion will contain vulnerabilities. Instead of attempting to learn bug patterns from this sparse signal, we advocate for the inverse. Since bugs are inherently anomalous, we believe that LLMs should be leveraged as a model for correct code. If we accept that code LLMs are capable of generalizing the structure of bug-free code, then bug detection can be cast as an anomaly detection problem.
@@
There have been previous attempts at modelling software defects and vulnerabilities as anomalous code. For example, studies of code “naturalness” [15, 16] have shown that source code contains many of the same characteristics as natural language and hence can be modelled by statistical methods. Subsequent work has then found a statistically significant difference between buggy and fixed code when measuring the entropy of predictions from a $n$-gram model trained on 35 million LOC, which enabled their approach to achieve bug detection that was competitive with static-analysis based tools.
@@
However, modern transformer-based LLMs have considerably greater representational power than $n$-gram models, and are trained on data sets with 10’s-100’s of billions of LOC (a 4-5 order of magnitude increase!). We hypothesize that such models will be far more capable of exploiting the insight that vulnerabilities are just anomalous code. But, modern LLMs also introduce a number of new challenges when applying them to anomaly-based vulnerability detection. First, modern LLMs are so accurate at predicting code, that they often predict correct code almost perfectly. In such cases, an entropy measure may underestimate the level of an anomaly between a perfect and an almost-perfect prediction, motivating an exploration of methods for measuring the anomaly level of predictions. Second, unlike the previous generation of $n$-gram models, LLMs can use a context of hundreds to thousands of tokens to make predictions, raising the question of the appropriate context size to use. Third, given the availability of a variety of industrially-trained open-source models of varying sizes and architectures, we find ourselves in a position to explore whether this approach truly generalizes to different model architectures, and how it performs as models scale to larger capacities and training sets. Finally, given that modern LLMs are trained on large training sets of code, there is a risk that a pre-trained LLM may have seen both the vulnerable and fixed version of some code – commonly referred to as “leakage”. We evaluate our approach on a leakage-free dataset we created that only contains vulnerabilities found after the training date of the LLMs we use to demonstrate that our results generalize to unseen vulnerabilities.
@@
  • We evaluate different parameters for our masked span reconstruction technique, namely the model size and prompt context. From our experiments, larger models and compound statement contexts enable more accurate vulnerability detection.
  • Our anomaly-based tool performs favourably with LineVul [8] and LineVD [9], two detectors that were trained to predict line-level vulnerabilities.
We make all code for our tool ANVIL available at https://anonymous.4open.science/r/anvil. All data and scripts used for running our evaluation are also made public for reproducibility.
@@ II-A Overview
Similar to previous research in code naturalness, we consider a span of code anomalous (i.e. unnatural) if a model considers it out-of-distribution. Specifically, the feedback we utilize is the LLM reconstruction accuracy on a code generation task, when masking each span of code for inference. LLMs are trained to predict the most probable token given a chain of previous tokens, hence a low reconstruction accuracy compared to the ground truth indicates that the ground truth has deviated from the learned representation of the model. Figure 1 describes the workflow of our technique, ANVIL. For each line of code under analysis, we withhold the original line as the ground truth and instruct the LLM to reconstruct it based on the surrounding context. The LLM generated code is then compared with the ground truth to compute an anomaly score, which quantifies the divergence between them.
@@ II-B FIM & Context Selection
Code LLMs are trained to perform specific generative tasks, such as code-completion, where the model is trained to predict the next $n$ tokens given a code prefix, or fill-in-the-middle (FIM, also known as infilling) [17], where the model is trained to generate tokens that fall between a given prefix and suffix.
@@
We focus on the utility of FIM as the mode of code generation for detecting anomalies. Although code completion is a more prevalent task, research by Bavarian et al. [17] highlights that FIM tasks generally result in lower reconstruction loss due to additional constraints provided by the suffix. This structure allows FIM to capture downstream control and data flows, which are crucial for generating correct code for a masked line.
@@
However, adapting the FIM task for anomaly detection introduces several challenges due to constraints such as GPU memory and LLM token limits, which make it impractical to use entire source files as context. To address this, we empirically evaluate two methods for finding a context size: (1) a fixed size approach and (2) a semantics-aware approach. For (1), we investigate the effect of fixing the context to a specific number of source-code lines, equally split between the prefix and suffix. For instance, a 500-line context means using 250 lines preceding and succeeding the line under analysis.
@@
Alternatively, (2) adaptively selects the maximum compound statement (MCS) surrounding the line under analysis. In C/C++, a compound statement refers to a sequence of statements enclosed by curly braces (“{” and “}”). Given a line of source code, we define the MCS as the largest compound statement that encapsulates the line. For each line, ANVIL statically extracts the MCS to use as the context. Often, the MCS equates to the function body containing the line; however, if the MCS is too large, exceeding token limits or GPU memory capacity, we resort to the next largest compound statement that meets these constraints. We note that this technique is agnostic to the hardware used, but empirically, we found that a context limit of 500 lines reached near-maximum peak usage of our GPU memory. Hence, our adaptive MCS selection only considers compound statement that are 500 lines or fewer.
@@
Figure 2: An example from libpng showing the largest and second-largest compound statement for a line of interest
@@
If no suitable compound statement is identified, we default to providing a prefix and suffix of 150 lines each. This choice is based on empirical data from our experiments, where a context window of 300 lines, among all fixed context sizes, struck the best balance between GPU memory utilization and reconstruction accuracy, leading to the most efficient differentiation between vulnerable and non-vulnerable lines. Additionally, when an MCS constitutes a function body, we include the function definition located above it to enrich the context.
@@
Figure 2 illustrates an example of how we extract the compound statements associated with a masked line. We use a stack to store the line numbers of all opening braces, and then match each closing brace with the latest unmatched opening brace on the top of the stack.
@@ II-C Anomaly Score
To quantify the discrepancies between the generated line and ground truth, we assign an anomaly score to each line of source code. We define the anomaly score as a function which takes the masked ground truth ($p$) and the generated line ($q$) as arguments and calculates a value $\delta(p,q)$, which estimates how “anomalous” the ground truth is. A higher score suggests a greater likelihood of the current line being a vulnerability. To concretize $\delta(p,q)$, an approach used in previous studies on code naturalness [16], is to compute the cross-entropy loss for each token generated by the LLM against the masked ground truth token. By averaging the loss values across the line, we obtain an overall loss measure ($\delta_{1}(p,q)=loss(p,q)$). While, this loss value ranges from 0 to infinity, in practice the loss typically ranges between 0 and 1.
An alternative metric to the anomaly score is using an exact match indicator, which assigns a value of 1 if the LLM’s generated line matches the ground truth exactly, and 0 otherwise. An exact match implies that, during model inference for each token, the most probable token in each token distribution is perfectly aligned with the ground truth, and hence the line under analysis should not be considered anomalous. Thus, we can use the negation of exact string match for anomaly scoring, $\delta_{2}(p,q)=-\mathbb{1}_{exact\_match}(p,q)$.
Using either metric alone may not fully capture the nuances of anomaly detection. The exact match indicator, being binary, produces an overly sparse distribution that may penalize slight deviations between the prediction and ground truth too harshly. The loss value, although continuous, might miss subtle yet critical discrepancies. On the one hand, it can under-estimate the level of anomaly: if a masked ground truth is “$if(a>=0)$” and the model predicts that the line is “$if(a>0)$”, the off-by-one-character ’=’ might result in a low average loss but represents a significant logical change, such as becoming the root cause of an out-of-bound array access. On the other hand, it can also over-estimate the level of anomaly: if the model has low certainty about many of the tokens in an otherwise perfectly correct prediction, it can end up computing a larger loss than the off-by-one character prediction above.
@@
$\delta(p,q)=loss(p,q)-\mathbb{1}_{exact\_match}(p,q)$ (1)
This hybrid approach combines both the loss and the exact match indicator, which allows us to heavily reward exact matches, signifying non-anomalous outputs; such matches typically result in a small anomaly score near -1. In comparison, non-matches, indicating anomalies, tend to result in higher scores larger than 0. However, because LLMs are prone to errors, even an exact match might still contain small possibilities that the line is vulnerable. Thus, we rely on the loss measure to capture finer discrepancies that could indicate vulnerabilities. In Section III-B, we have conducted experiments comparing all three functions, and validated that the hybrid approach offers the most reliable performance in identifying vulnerabilities as anomalous.

### v2

@@ Abstract
Supervised-learning-based software vulnerability detectors often fall short due to the inadequate availability of labelled training data. In contrast, Large Language Models (LLMs) such as GPT-4, are not trained on labelled data, but when prompted to detect vulnerabilities, LLM prediction accuracy is only marginally better than random guessing. In this paper, we explore a different approach by reframing vulnerability detection with LLMs as anomaly detection. Since the vast majority of code does not contain vulnerabilities and LLMs are trained on massive amounts of such code, vulnerable code can be viewed as an anomaly when compared to LLM’s predicted code, freeing the model from the need for labelled data to provide a learnable representation of vulnerable code. Leveraging this perspective, we used LLMs to perform a masked code reconstruction task, and compared the LLM-generated code to its original version to identify anomalies. Our experiments demonstrate that LLMs trained for code generation exhibit a significant gap in reconstruction accuracy between vulnerable and non-vulnerable code, which can be quantified using an anomaly score.
Using this insight, we implement ANVIL, a detector that identifies software vulnerabilities at line-level granularity. Our experiments explore the discriminating power of different anomaly scoring methods, as well as the sensitivity of ANVIL to context size. We also study the effectiveness of ANVIL on various LLM families, and conduct leakage experiments on vulnerabilities that were discovered after the knowledge cutoff of our evaluated LLMs. On a collection of vulnerabilities from the PrimeVul dataset, ANVIL outperforms state-of-the-art line-level vulnerability detectors, LineVul, LineVD, and LLMAO, which have been trained with labelled data, despite ANVIL having never been trained with labelled vulnerabilities. Specifically, our approach achieves 1.25×1.25$\times$1.25 × to 1.62×1.62$\times$1.62 × better Top-5 accuracies, 27% to 69% better Normalized MFR, and 1.10×1.10$\times$1.10 × to 1.15×1.15$\times$1.15 × times better ROC scores on line-level vulnerability detection tasks.
@@ I Introduction
Probabilistic approaches to static vulnerability detection have garnered considerable research attention due to the promise of automatically detecting common bug patterns by training a model on historical data. Confidence in statistical modelling has only grown with the recent advances in large language models (LLMs). One of the significant advantages of LLMs is that they can be trained in a self-supervised fashion without labelled training data. This has enabled them to be trained on very large training sets. This approach has led to impressive and surprising results in code understanding and code generation [feng_codebert_2020, wang_codet5_2021].
@@
Unfortunately, the same cannot be said about their ability to detect vulnerabilities. As shown in a recent study [steenhoek_comprehensive_2024], even with cutting edge LLMs like GPT-4, vulnerability prediction accuracy on real world code is little better than a coin-flip. One reason for this limitation is that the pretraining process of LLMs is designed primarily to generate correct code, without explicitly incorporating the concept of “vulnerability”, leaving them with insufficient domain knowledge to reason about whether a given code is vulnerable. On the other hand, current learning and LLM-based vulnerability detectors, while attempting to embed the notion of code vulnerabilities, still suffer from the lack of a large and well-labelled set of representative vulnerability dataset [chen_diversevul_2023, ding_vulnerability_2024]. Efforts to compensate for the dearth of labelled data by selecting relevant code features (e.g., the granularity of the code context, commit history, control-flow, data-flow, etc.) and converting these features into a learnable representation (e.g., a vector or a graph) [li_vuldeepecker_2018, li_vulnerability_2021, cheng_how_2022, fu_linevul_2022, hin_linevd_2022, mirsky_vulchecker_2023, russell_automated_2018] have only yielded modest improvements. Similarly, recent attempts to generate new vulnerable training corpora [mirsky_vulchecker_2023, nong_vulgen_2023, nong_vgx_2024] have not improved real-world performance significantly, as synthetically generated vulnerabilities still cannot capture the huge variety of ways that vulnerabilities can manifest themselves.
@@
LLMs are effective because they are trained on massive datasets, enabling them to learn the general structure and patterns required to generate correct code. However, this broad understanding does not translate well to recognizing vulnerable code patterns due to fundamentally imbalanced corpora. Out of the huge amounts of unlabelled code that LLMs are trained on, only a minuscule portion will contain vulnerabilities. Instead of attempting to learn bug patterns from this sparse signal, we advocate for the inverse. Since bugs are inherently anomalous, we believe that LLMs should be leveraged as a model for correct code, and bug detection can be cast as an anomaly detection problem. By comparing the original code to the LLM-generated version in a masked code reconstruction task, we can determine whether the original code falls outside the LLM’s generation distribution, thereby identifying anomalies.
{comment}
@@
There have been previous attempts at modelling software defects and vulnerabilities as anomalous code. For example, studies of code “naturalness” [hindle_naturalness_2016] have shown that source code contains many of the same characteristics as natural language and hence can be modelled by statistical methods. Subsequent work[ray_naturalness_2016] used a purely statistical $10$-gram model trained on 35 million LOC for next-token prediction and found a statistically significant difference in entropy (a measure of prediction uncertainty) when predicting buggy versus fixed code, which enabled their approach to achieve bug detection that was competitive with static-analysis-based tools. There have been previous attempts at modelling software defects and vulnerabilities as anomalous code. For example, a study of code “naturalness” [ray_naturalness_2016] leveraged statistical methods to model software bugs and their fixes, creating a purely statistical $10$-gram model trained on 35 million LOC for next-token prediction. Then, given a sequence (i.e., line of code), the model returned a probability of that sequence based on its internal statistics. Using this probability, the approach calculated the cross-entropy and using this data, the work identified a statistically significant difference in entropy (a measure of prediction uncertainty) when evaluating buggy versus fixed code, which enabled their approach to achieve bug detection that was competitive with static-analysis-based tools.
{comment}
@@
However, modern transformer-based LLMs have considerably greater representational power than $n$-gram models, and are trained on data sets with 10’s-100’s of billions of LOC (a 4-5 order of magnitude increase). We hypothesize that such models will be far more capable of exploiting the insight that vulnerabilities are just anomalous code. However, modern LLMs also introduce a number of new challenges when applied to anomaly-based vulnerability detection. First, LLMs are very accurate at predicting correct code and, thus, have the potential to detect subtle vulnerabilities, even when the difference between vulnerable code and its patched version is as small as a single character. For example, an out-of-bound array access may result from a missing equal sign in an “if” statement. In such cases, the predicted correct code closely resembles the original vulnerable version, making the entropy difference between the two negligible. This poses a challenge, as small entropy differences make it difficult to establish a meaningful threshold for detecting vulnerabilities effectively. To address this limitation, this paper explores different methods for quantifying the anomaly level of predictions.color=blue!60color=blue!60todo: color=blue!60Joe: Hopefully it improves the flow Second, unlike the previous generation of $n$-gram models, LLMs can use a context of hundreds to thousands of tokens to make predictions, raising the question of the appropriate context size to use. Third, given the availability of a variety of industrially-trained open-source models of varying sizes and architectures, we find ourselves in a position to explore whether this approach truly generalizes to different model architectures, and how it performs as models scale to larger capacities and training sets. Finally, given that modern LLMs are trained on large training sets of code, there is a risk that a pre-trained LLM may have seen both the vulnerable and fixed version of some code – commonly referred to as “leakage”. We evaluate our approach on a leakage-free dataset we created that only contains vulnerabilities found after the training date of the LLMs we use to demonstrate that our results generalize to unseen vulnerabilities.
We hypothesize that the 10-gram model result can be improved by the modern transformer-based LLMs since they have considerably greater representational power than $n$-gram models, being trained on data sets with 10’s-100’s of billions of LOC (a 4-5 order of magnitude increase). However, the application of LLMs to anomaly-based vulnerability detection poses new challenges. First, measuring the entropy of predictions might be too coarse when applied to LLM results because LLMs are so precise at generating code that even a one character difference can be significant.color=purple!30color=purple!30todo: color=purple!30david For example, an out-of-bound array access may result from a missing equal sign in an “if” statement. To address this limitation, this paper explores different methods for quantifying the anomaly level of predictions. Second, unlike the previous generation of $10$-gram models, LLMs can use a context of hundreds to thousands of tokens to make predictions, raising the question of the appropriate context size to use. Third, the design space of LLMs is very large, with a diversity of model architectures, training objectives and datasets. Thus, we also explore whether anomaly-based bug detection with LLMs generalizes to different model architectures and how it performs as models scale to larger capacities and training sets. Finally, given that modern LLMs are trained on large training sets of code, there is a risk that a pre-trained LLM may have seen both the vulnerable and fixed version of some code – commonly referred to as “leakage”. We evaluate our approach on a leakage-free dataset we created that only contains vulnerabilities found after the training date of the LLMs we use to demonstrate that our results generalize to unseen vulnerabilities.
@@
  • We evaluate different hyperparameters for our masked code reconstruction technique, namely the anomaly scoring function, model size and prompt context. From our experiments, the hybrid anomaly function, larger models and compound statement contexts enable more accurate and efficient vulnerability detection.
  • Our anomaly-based tool outperforms LineVul [fu_linevul_2022], LineVD [hin_linevd_2022], and LLMAO [yang_large_2024], three state-of-the-art LLM-based detectors trained to predict line-level vulnerabilities. Specifically, ANVIL achieves 1.25× to 1.62× better Top-5 accuracies and 27% to 69% improvements in Normalized MFR for vulnerability prioritization, as well as 1.10× to 1.15× higher ROC-AUC scores for line-level vulnerability classification.
@@ II-A Overview
Similar to previous research [ray_naturalness_2016] in code “naturalness”, we consider a piece of code anomalous (i.e., unnatural) if a model considers it out-of-distribution. Specifically, the feedback we utilize is the LLM reconstruction accuracy on a code generation task, when masking each line of code for inference. LLMs are trained to predict the most probable token given a chain of previous tokens, hence a low reconstruction accuracy compared to the ground truth indicates that the ground truth has deviated from the learned representation of the model. Figure 1 describes the workflow of our technique, ANVIL. For each line of code under analysis, we withhold the original line as the ground truth and instruct the LLM to reconstruct it based on the surrounding context. The LLM-generated code is then compared with the ground truth to compute an anomaly score, which quantifies the divergence between them. While in our study, we only implement ANVIL to support C/C++ languages as a proof of concept, we believe the underlying approach can be generalized to detect vulnerabilities in other programming languages.
@@ II-B FIM & Context Selection
Code LLMs are trained to perform specific generative tasks, such as code-completion, where the model is trained to predict the next $n$ tokens given a code prefix, or fill-in-the-middle (FIM, also known as infilling) [bavarian_efficient_2022], where the model is trained to generate tokens that fall between a given prefix and suffix.
@@
We focus on the utility of FIM as the mode of code generation for detecting anomalies. Although code completion is a more prevalent task, research by Bavarian et al. [bavarian_efficient_2022] highlights that FIM tasks generally result in lower reconstruction loss due to additional constraints provided by the suffix. This structure allows FIM to capture downstream control and data flows, which are crucial for generating correct code for a masked line.
@@
However, adapting the FIM task for anomaly detection introduces several challenges due to constraints such as GPU memory and LLM token limits, which make it impractical to use entire source files as context. To address this, we empirically evaluate two methods for finding a context size: (1) a fixed size approach and (2) a semantics-aware approach.
For (1), we investigate the effect of fixing the context to a specific number of source-code lines, equally split between the prefix and suffix. For instance, a 500-line context means using 250 lines preceding and succeeding the line under analysis. Still, a fixed context length, such as a 150-line prefix and suffix, may not sufficiently encapsulate all relevant code features for vulnerability detection. In cases where the relevant context extends beyond this 300-line window, such as in functions longer than 300 lines, critical contextual information will not be conveyed to the LLM. Conversely, in smaller functions that span only 50 lines, a 300-line context window could introduce irrelevant information.
@@
Alternatively, (2) uses an Adaptive Context (AC) approach, which adaptively selects the maximum compound statement surrounding the line under analysis. This method ensures that the model receives semantically relevant context to effectively learn the code’s features without being burdened by irrelevant information that the model might inherently ignore. In C/C++, a compound statement refers to a sequence of statements enclosed by curly braces (“{” and “}”). Given a line of source code, we define the AC as the largest compound statement that encapsulates the line. For each line, ANVIL statically extracts the AC to use as the context. Often, the AC equates to the function body containing the line; however, if the AC is too large, exceeding token limits or GPU memory capacity, we resort to the next smaller compound statement that meets these constraints. We note that this technique is agnostic to the hardware used, but empirically, we found that a context limit of 500 lines reached near-maximum peak usage of our GPU memory. Hence, our AC selection only considers compound statements that are 500 lines or fewer.
@@
Figure 2: An example from lua (a scripting language developed in C) showing the largest and second-largest compound statement for a line of interest
@@
If no suitable compound statement is identified, we default to providing a prefix and suffix of 150 lines each. This choice is based on empirical data from our experiments, where a context window of 300 lines, among all fixed context sizes, struck the best balance between GPU memory utilization and reconstruction accuracy, leading to the most efficient differentiation between vulnerable and non-vulnerable lines.
@@
Figure 2 illustrates an example of how we extract the compound statements associated with a masked line. Since our analysis is performed at the source code level, we rely on matching brackets to identify compound statements. Specifically, we use a stack to store the line numbers of all opening braces, and then match each closing brace with the latest unmatched opening brace on the top of the stack.
@@ II-C Anomaly Score
To quantify the discrepancies between the generated line and ground truth, we assign an anomaly score to each generated line. We define the anomaly score as a function that takes the masked ground truth ($p$) and the generated line ($q$) as arguments and calculates a value $\delta(p,q)$, which estimates how “anomalous” the ground truth is. A higher score suggests a greater likelihood of the current line being a vulnerability.
Previous work on code “naturalness” [ray_naturalness_2016] used a similar metric to this anomaly score, which calculates a cross-entropy loss for each token based on $10$-gram frequencies counted from their corpus. In our study, we compute the anomaly score as the average cross-entropy loss across the entire reconstructed line generated by the LLM compared to the masked ground truth. This yields an overall loss measure $\delta_{loss}(p,q)=loss(p,q)$ for the line. Although the loss theoretically ranges from 0 to infinity, in practice, we observed that it typically falls between 0 and 1 in our experiments. color=blue!60color=blue!60todo: color=blue!60joe
{comment}
To concretize $\delta(p,q)$, an approach used in previous studies on code naturalness [ray_naturalness_2016], is to compute the cross-entropy loss for each token generated by the LLM against the masked ground truth token. By averaging the loss values across the line, we obtain an overall loss measure ($\delta_{loss}(p,q)=loss(p,q)$). While, this loss value ranges from 0 to infinity, in our experiments, we found that the loss typically falls between 0 and 1.
Alternatively, we propose using an exact match indicator as another metric. This approach assigns a value of 1 if the LLM’s generated line matches the ground truth exactly, and 0 otherwise. An exact match implies that, during model inference for each token, the most probable token in each token distribution is perfectly aligned with the ground truth, and hence the line under analysis should not be considered anomalous. Thus, we can use the negation of exact string match for anomaly scoring, $\delta_{EM}(p,q)=-\mathbb{1}_{exact\_match}(p,q)$.
As LLMs are very accurate, using either metric alone may not fully capture the nuances of using LLMs for anomaly detection. The exact match indicator, being binary, produces an overly sparse distribution that may penalize slight deviations between the prediction and ground truth too harshly. The loss value, although continuous, might miss subtle yet critical discrepancies. On the one hand, it can under-estimate the level of anomaly: if a masked ground truth is “$if(a>=0)$” and the model predicts that the line is “$if(a>0)$”, the off-by-one-character ’=’ might result in a low average loss but represents a significant logical change, such as becoming the root cause of an out-of-bound array access. On the other hand, it can also over-estimate the level of anomaly: if the model has low certainty about many of the tokens in an otherwise perfectly correct prediction, it can end up computing a larger loss than the off-by-one character prediction above.
@@
$\delta_{Hybrid}(p,q)=loss(p,q)-\mathbb{1}_{exact\_match}(p,q)$ (1)
This hybrid approach combines both the loss and the exact match indicator, which allows us to heavily reward exact matches, signifying non-anomalous outputs; such matches typically result in a small anomaly score near -1. In comparison, non-matches, indicating anomalies, tend to result in higher scores larger than 0. However, if LLM generates code with an error, the exact match would say completely the opposite, since we entrust LLMs to generate bug-free code. Thus, we rely on the loss measure to capture finer discrepancies that could still indicate vulnerabilities. In Section III-B, we have conducted experiments comparing all three functions, and validated that the hybrid approach offers the most reliable performance in identifying vulnerabilities as anomalous.